NEWER(1) NEWER(1) NAME newer−comparefilemodificationtimes SYNOPSIS newer[option]file1file2 DESCRIPTION Exitsuccessfullyiffile1existsandisatleastasoldasfile2. OPTIONS neweracceptsthefollowingoptions: --helpPrinthelpmessageandexit. --version Printversioninformationandexit. Web2C7.32x 4January1998 1